OpenStudy (anonymous):

How do I figure out w/4-4=3?

OpenStudy (anonymous):

Add \(4\) both the sides first..

OpenStudy (anonymous):

\[\frac{w}{4} -4 + \color{red}{4} = 3 + \color{red}{4}\]

OpenStudy (anonymous):

On left hand side, -4 and +4 cancels out.. \[\frac{w}{4} = 7\]

OpenStudy (anonymous):

Now, multiply by \(4\) both the sides..

OpenStudy (anonymous):

\[\frac{w}{4} \times \color{blue}{4} = 7 \times \color{blue}{4} \\ w = 7 \times 4 = ??\]

OpenStudy (anonymous):

So it would be 28?

OpenStudy (anonymous):

Yes, you want to check?? then you can check too.. Put \(w = 28\) in left side, you will get right side..

OpenStudy (anonymous):